# Advancing the Database of Cross-Linguistic Colexifications with New Workflows and Data

### v1

@@ Abstract
Lexical resources are crucial for cross-linguistic analysis and can provide new insights into computational models for natural language learning. Here, we present an advanced database for comparative studies of words with multiple meanings, a phenomenon known as colexification. The new version includes improvements in the handling, selection and presentation of the data. We compare the new database with previous versions and find that our improvements provide a more balanced sample covering more language families worldwide, with an enhanced data quality, given that all word forms are provided in phonetic transcription. We conclude that the new Database of Cross-Linguistic Colexifications has the potential to inspire exciting new studies that link cross-linguistic data to open questions in linguistic typology, historical linguistics, psycholinguistics, and computational linguistics.
@@ 1 Introduction
The Database of Cross-Linguistic Colexifications (CLICS, https://clics.clld.org, Rzymski et al., 2020) offers detailed data on the distribution and frequency of colexifications across several thousand languages. Colexification is a cover term that unifies the notion of polysemy, homophony, and underspecification, referring to cases where a single word form in a given language expresses multiple senses (François, 2008). For example, Vietnamese xanh refers to ‘blue’ and ‘green’ at the same time, German böse means both ‘angry’ and ‘evil’, or English ear refers to a part of the body or a part of a grain. The different examples represent words with multiple senses and can be labelled as underspecification (Vietnamese), polysemy (German), or homophony (English), but they can also be taken together as examples of the phenomenon of colexification.
CLICS has built on this idea by collecting data from multilingual word lists that were unified with respect to the semantic glosses by which words across different languages are elicited. From these word lists, colexifications were automatically extracted, forming a large colexification network (List et al., 2013) that can be investigated interactively (Mayer et al., 2014). The database has improved concerning the workflow by which data are aggregated and in terms of the number of datasets underlying the database (4 datasets in Version 1.0, List et al. 2014, 15 datasets in Version 2.0, List et al. 2018, 30 datasets in Version 3.0 Rzymski et al. 2020). In its current form, the CLICS database is characterized by three major features. First, CLICS aggregates data from existing standardized datasets, rather than curating data directly. Second, CLICS offers its data in both machine- and human-readable form, allowing scholars to access the data in computational workflows as well as through the web interface. Third, CLICS is open, and both the individual data and the source code are published with permissive licenses, allowing scholars not only to investigate the database, but also to extend it, or to provide additional content on top of it.
@@
Given that five years have passed since the last official release of CLICS and that new relevant datasets have been published during this time, mainly as part of Lexibank, a large repository for standardized multilingual word lists (https://lexibank.clld.org, List et al., 2022; Blum et al., 2025), it is time to improve the database even further. Taking advantage of the fact that CLICS is open and free to modification, we therefore present an updated version of the CLICS database, which we named CLICS* for convenience. CLICS* not only increases the underlying data, but also addresses three major shortcomings of the previous versions of CLICS by improving (1) the handling of concepts (§ 3.2), (2) the selection of languages to be included in the colexification database (§ 3.3), and (3) the general representation of data (§ 3.4). In the following, we will present previous studies devoted to cross-linguistic colexifications (§ 2.1), discuss the improvements in more detail (§ 3), illustrate their consequences for CLICS* (§ 4), and reflect on the future of cross-linguistic colexification data (§ 5).
@@ 2.1 Cross-Linguistic Colexifications
Not long after François (2008) had first introduced the term colexification along with initial ideas how the phenomenon could be analyzed using cross-linguistic data, typologists quickly adopted the term and the technique to study lexical semantics both globally and in certain linguistic areas. Two major reasons contributed to the popularity of the term and the technique.
First, polysemy and homophony are notoriously difficult to distinguish, specifically when analysing languages whose history is less well known. While scholars sometimes distinguish both relations by degree of semantic similarity, arguing that homophonous words show greater divergence in meaning than polysemous words (Leivada and Murphy, 2021, 7), the original distinction between polysemy and homophony is strictly diachronic. Thus, they reflect two distinct processes of language change: polysemy is the result of semantic change, while homophony is the result of a merger of originally distinct word forms due to sound change (Sperber 1923, 12f, Apresjan 1974, 11). However, in the mind of speakers, the history of the words does not play any major role. Speakers seem to show some general awareness that some words have multiple senses that are closely related to each other, whereas other words with distinct senses merely sound alike (Enfield and Comrie, 2015, 20f). While it may seem useful to distinguish polysemy and homophony in theory, the distinction of the two relations in practice is difficult to make. Omitting the explicit distinction between the two forms of lexical ambiguity allowed scholars to assemble data in an unbiased and efficient way. Scholars could accumulate colexification data for their areas of interest without having to discuss the consequences of unpractical terminology. Instead of deciding whether the findings would reflect polysemy or homophony, scholars could let the data decide, given that polysemy often largely exceeds homophony.
Second, scholars began to explore the benefits of modelling cross-linguistic colexifications with the help of network approaches (Cysouw, 2010). This not only led to clear visualizations of semantic similarities that could be observed across languages, but also opened up new possibilities for the analysis of cross-linguistic polysemy using network approaches (List et al., 2013) and the introduction of interactive techniques for data visualization and exploration, which later became a core component of the CLICS database (Mayer et al., 2014). Figure 1 illustrates how colexification networks can be constructed from colexification data, using data from CLICS 3 (Rzymski et al., 2020).
@@ 2.2 Data Aggregation and Analysis in CLICS
The integral part of the Database of Cross-Linguistic Colexifications is the workflow by which data are aggregated from individual datasets and later analyzed to create a colexification network. To be able to aggregate data from different sources, source datasets must be standardized. Standardization is achieved with the help of Cross-Linguistic Data Formats (CLDF, https://cldf.clld.org, Forkel et al., 2018), an initiative that builds on the CSVW standard for tabular data on the web (https://csvw.org, Gower, 2021), but extends CSVW with semantics relevant to comparative linguistics. A CLDF dataset is a collection of CSV files linked via a JSON file that stores the metadata, providing information on how the CSV files should be interpreted computationally and what values are shared across the files. Thus, a CLDF dataset is a small relational database with specific semantics that link the data with additional data from outside.
@@
The most important external datasets that CLICS links to are three reference catalogs: Glottolog, Concepticon, and CLTS. Glottolog (https://glottolog.org, Hammarström et al., 2024) offers basic information on language varieties, including information on language classification, geolocations, and the documentation status of individual languages. Concepticon (https://concepticon.clld.org, List et al., 2025) offers a collection of basic senses that are expressed across multilingual word lists. Senses are provided in the form of concept sets that are linked across several hundred concept lists that have been annotated by the Concepticon team in the past decade (for details on the curation process, see Tjuka et al., 2023). CLTS (https://clts.clld.org, List et al., 2021) is a reference catalog for Cross-Linguistic Transcription Systems that standardizes phonetic transcriptions by advocating a subset of the International Phonetic Alphabet (IPA, International Phonetic Association, 1999) that is represented in the form of distinctive features (for details, see Anderson et al. 2018 and Rubehn et al. 2024). The conversion of individual datasets to the CLDF standard is supported by dedicated Python libraries (most importantly the CLDFBench packages, Forkel and List 2020) that help to check the overall consistency of the data.
@@
From a collection of CLDF datasets, the CLICS aggregation workflow iterates over the datasets and assembles cross-linguistic colexifications for each language variety. Here, CLICS uses an efficient method that avoids comparing $n$ words in one language against $n$ words in the same languages, but rather identifies colexifications from tuples, consisting of a word form and its corresponding sense (a Concept Set in the Concepticon catalog), with the help of hash tables (List, 2022).
@@
Having created a large colexification network of all CLDF datasets, the CLICS workflow analyzes this data further, by computing communities, that is, partitions of nodes in a graph that show more connections to each other than to other nodes outside of the partition (Newman, 2006, 8577). Communities are inferred with the help of the Infomap algorithm (Rosvall and Bergstrom, 2008) and are used to structure the web application, by allowing users to inspect either entire communities or individual subsets of the data. The methods for data aggregation and analysis are freely accessible and can be easily applied by scholars to create their analyses of subsets of the data in CLICS or by extending the CLICS collection further, as illustrated, for example, by Tjuka (2024b).
@@ 2.3 Shortcomings of CLICS
The first shortcoming relates to the data underlying CLICS. While data from 30 datasets were aggregated in Version 3 (Rzymski et al., 2020), many more datasets have recently been made available via the Lexibank repository (List et al., 2022). Improving the database by increasing the number of datasets is thus one of the most urgent tasks that should be addressed in an updated version.
@@
Third, CLICS 3 provided information from more than 3,000 language varieties. However, a closer look at the data showed that only a small proportion of the included languages met the requirement set by the editors of CLICS 3 to provide translations for at least 250 concepts. For CLICS 3, the authors instead selected 30 data sets that were officially compiled from concept lists with 250 or more elements. The resulting word lists for individual languages, however, were often scarce and a larger number of the languages did not meet the originally stated coverage criterion.
Fourth, CLICS 3 offered the colexification network exclusively in the form of a GML-file. Although GML is a a common format for the encoding of graphs (Himsolt, 2010), accepted by many software tools, including igraph (https://igraph.org, Csárdi and Nepusz, 2006), NetworkX (https://networkx.org/, Hagberg et al., 2008), and Cytoscape (http://cytoscape.org/, Smoot et al., 2011), the format is not well-suited to share the extensive data on colexification patterns computed by CLICS 3. As a result, more transparent data formats for handling colexification data and colexification networks are needed to represent the results of the CLICS workflow in due detail.
@@ 3 Materials and Methods
In the following, we will introduce all necessary steps that lead to the creation of our modified CLICS* database. We followed the established workflow for data aggregation used in CLICS 3 to some extent (§ 2.2). However, we present a drastic increase of data based on standardized datasets (§ 3.1), introduce an improved handling of concepts during data aggregation (§ 3.2), refine the selection of languages and concepts (§ 3.3), and make the representation of the colexification data more transparent (§ 3.4).
@@ 3.1 Data Basis
CLICS 3 was based on 30 datasets available in CLDF. Many more datasets have since been published as part of the Lexibank repository, which was first published in 2022 (List et al., 2022) as Lexibank 1.0 and curates data from 100 different datasets of different sizes. Of those 52 Lexibank datasets were suitable for the inclusion in CLICS, because they were based on concept lists that contain 250 and more items (this criterion was used to build CLICS 3, Rzymski et al. 2020). The newest version, Lexibank 2.0, offers data for 134 different datasets that are all phonetically transcribed (Blum et al., 2025). For our enhanced version of CLICS, we identified 94 suitable datasets. These datasets are listed in the supplementary material accompanying this study.
@@ 3.2 Concept Handling
The colexifications in CLICS result from comparing words mapped to the standardized concept sets in Concepticon (List et al., 2016; Tjuka et al., 2023). The consequent mapping of the elicitation glosses in individual datasets to the Concepticon reference catalog has been one of the most important factors that allowed for the growth of CLICS: Version 1 (List et al., 2014) containing 221 language varieties and 1,280 concepts, Version 2 (List et al., 2018) containing 1,220 language varieties and 2,487 concepts, and Version 3 (Rzymski et al., 2020) containing 3,156 language varieties and 2,906 concepts. However, through the mapping of the datasets to the Concepticon, a bias for a certain number of concepts that exhibit hierarchical relations to other concepts was introduced.
@@
Already with its first launch (List et al., 2016), Conception has allowed for the definition of broad concepts that are expressed as such only in specific languages or specific linguistic areas. As an example, consider the concept sets arm or hand and foot or leg. These concepts are expressed by individual word forms in languages such as Vietnamese tay, referring to ‘arm’ or ‘hand’, or Russian noga, referring to ‘foot’ or ‘leg’. However, many languages distinguish them further, using individual words for arm, hand, foot, and leg, respectively.
@@
Some lists in Concepticon have a linguistic area or language family as a target. Thus, the introduction of underspecified concept sets, such as ARM OR HAND or FOOT OR LEG was important, because linguists reporting on Slavic languages or particular languages in South-East Asia do not elicit both arm and hand, if they know that these are always colexified in the languages under study. However, this kind of lexical underspecification, as we encounter it in the lexicon of Vietnamese and Russian, is one of the typical reasons for colexifications. Therefore, it is important to list such cases as true colexifications of arm and hand, as well as foot and leg. The original aggregation technique used by CLICS ignores these cases. As a result, important colexification information for a larger number of languages is lost.
@@
In our updated version CLICS*, we account for underspecification directly, by defining a list of 85 concept sets that exhibit underspecification along with the more specific target concepts that they cover. While most of these underspecified concept sets can be represented by two concept sets, some are represented by more than two (specifically kinship terms like sister and brother, which have four counterparts: younger sister (of man), younger sister (of woman), older sister (of man), and older sister (of woman for the former). In addition, we decided to replace some concept sets with a too broad or too narrow definition by more common concept sets (e.g. replacing stone or rock by stone because rock did not occur in the data).
@@
When encountering words that are mapped to these concepts during the initial iteration over all word lists in the data, the respective words are multiplied and each of the words is mapped to the specific concept sets covered by the underspecified concept sets. Word forms that are artificially multiplied in this form are marked in the resulting dataset by providing information on the original concept set. In total, we identify 85 underspecified concept sets in Concepticon that are relevant for the data in our modified version of CLICS. Of the 1,443,325 word forms in CLICS*, 107 921 word forms result from this refinement procedure. A detailed list of the concept replacements can be found in Appendix A.

### v2

@@ Abstract
Lexical resources are crucial for cross-linguistic analysis and can provide new insights into computational models for natural language learning. Here, we present an advanced database for comparative studies of words with multiple meanings, a phenomenon known as colexification. The new version includes improvements in the handling, selection and presentation of the data. We compare the new database with previous versions and find that our improvements provide a more balanced sample covering more language families worldwide, with enhanced data quality, given that all word forms are provided in phonetic transcription. We conclude that the new Database of Cross-Linguistic Colexifications has the potential to inspire exciting new studies that link cross-linguistic data to open questions in linguistic typology, historical linguistics, psycholinguistics, and computational linguistics.
@@ 1 Introduction
The Database of Cross-Linguistic Colexifications (CLICS, https://clics.clld.org, Rzymski et al., 2020) offers detailed data on the distribution and frequency of colexifications across several thousand languages. Colexification is a cover term that unifies the notions of polysemy, homophony, and underspecification, referring to cases where a single word form in a given language expresses multiple senses (François, 2008). For example, Vietnamese xanh refers to ‘blue’ and ‘green’ at the same time, German böse means both ‘angry’ and ‘evil’, or English ear refers to a part of the body or a part of a grain. The different examples represent words with multiple senses and can be labeled as underspecification (Vietnamese), polysemy (German), or homophony (English), but they can also be taken together as examples of the phenomenon of colexification.
CLICS has built on this idea by collecting data from multilingual word lists that were unified with respect to the semantic glosses by which words across different languages are elicited. From these word lists, colexifications were automatically extracted, forming a large colexification network (List et al., 2013) that can be investigated interactively (Mayer et al., 2014). The database has improved concerning the workflow by which data are aggregated and in terms of the number of datasets underlying the database (4 datasets in Version 1.0, List et al. 2014, 15 datasets in Version 2.0, List et al. 2018, 30 datasets in Version 3.0 Rzymski et al. 2020). In its current form, the CLICS database is characterized by three major features. First, CLICS aggregates data from existing standardized datasets, rather than curating data directly. Second, CLICS offers its data in both machine- and human-readable form, allowing scholars to access the data in computational workflows as well as through the web interface. Third, CLICS is open, and both the individual data and the source code are published with permissive licenses, allowing scholars not only to investigate the database, but also to extend it with additional content or methods.
@@
Given that five years have passed since the last official release of CLICS and that new relevant datasets have been published during this time, mainly as part of Lexibank, a large repository for standardized multilingual word lists (https://lexibank.clld.org, List et al., 2022; Blum et al., 2025), it is time to improve the database even further. Taking advantage of the fact that CLICS is open and free to modification, we therefore present an updated version of the CLICS database, which we named CLICS 4 for convenience. CLICS 4 not only increases the underlying data, but also addresses three major shortcomings of the previous versions of CLICS by improving (1) the handling of concepts (§ 3.2), (2) the selection of languages to be included in the colexification database (§ 3.3), and (3) the general representation of data (§ 3.4). In the following, we will present previous studies devoted to cross-linguistic colexifications (§ 2.1), discuss the improvements in more detail (§ 3), illustrate their consequences for CLICS 4 (§ 4), and reflect on the future of cross-linguistic colexification data (§ 5).
@@ 2.1 Cross-Linguistic Colexifications
Not long after François (2008) had first introduced the term colexification along with initial ideas on how the phenomenon could be analyzed using cross-linguistic data, typologists quickly adopted the term and the technique to study lexical semantics both globally and in certain linguistic areas. Two major reasons contributed to the popularity of the term and the technique.
First, polysemy and homophony are notoriously difficult to distinguish, specifically when analyzing languages whose history is less well known. While scholars sometimes distinguish both relations by degree of semantic similarity, arguing that homophonous words show greater divergence in meaning than polysemous words (Leivada and Murphy, 2021, 7), the original distinction between polysemy and homophony is strictly diachronic. Thus, they reflect two distinct processes of language change: polysemy is the result of semantic change, while homophony is the result of a merger of originally distinct word forms due to sound change (Sperber 1923, 12f, Apresjan 1974, 11). However, in the minds of speakers, the history of the words does not play a major role. Speakers seem to show some general awareness that some words have multiple senses that are closely related to each other, whereas other words with distinct senses merely sound alike (Enfield and Comrie, 2015, 20f). While it may seem useful to distinguish polysemy and homophony in theory, the distinction of the two relations in practice is difficult to make. Omitting the explicit distinction between the two forms of lexical ambiguity allowed scholars to assemble data in an unbiased and efficient way. Scholars could accumulate colexification data for their areas of interest without having to discuss the consequences of impractical terminology. Instead of deciding whether the findings would reflect polysemy or homophony, scholars could let the data decide, given that polysemy often largely exceeds homophony.
Second, scholars began to explore the benefits of modeling cross-linguistic colexifications with the help of network approaches (Cysouw, 2010). This not only led to clear visualizations of semantic similarities that could be observed across languages, but also opened up new possibilities for the analysis of cross-linguistic polysemy using network approaches (List et al., 2013) and the introduction of interactive techniques for data visualization and exploration, which later became a core component of the CLICS database (Mayer et al., 2014). Figure 1 illustrates how colexification networks can be constructed from colexification data, using data from CLICS 3 (Rzymski et al., 2020).
@@ 2.2 Data Aggregation and Analysis in CLICS
The integral part of the Database of Cross-Linguistic Colexifications is the workflow by which data are aggregated from individual datasets and later analyzed to create a colexification network. To be able to aggregate data from different resources, datasets must be standardized. Standardization is achieved with the help of Cross-Linguistic Data Formats (CLDF, https://cldf.clld.org, Forkel et al., 2018), an initiative that builds on the CSVW standard for tabular data on the web (https://csvw.org, Gower, 2021), but extends CSVW with semantics relevant to comparative linguistics. A CLDF dataset is a collection of CSV files linked via a JSON file that stores the metadata, providing information on how the CSV files should be interpreted computationally and what values are shared across the files. Thus, a CLDF dataset is a small relational database with specific semantics that link the data with additional data from outside.
@@
The most important external datasets that CLICS links to are three reference catalogs: Glottolog, Concepticon, and CLTS. Glottolog (https://glottolog.org, Hammarström et al., 2025) offers basic information on language varieties, including information on language classification, geolocations, and the documentation status of individual languages. Concepticon (https://concepticon.clld.org, List et al., 2025) offers a collection of basic senses that are expressed across multilingual word lists. Senses are provided in the form of concept sets that are linked across several hundred concept lists that have been annotated by the Concepticon team in the past decade (for details on the curation process, see Tjuka et al., 2023). CLTS (https://clts.clld.org, List et al., 2021) is a reference catalog for Cross-Linguistic Transcription Systems that standardizes phonetic transcriptions by advocating a subset of the International Phonetic Alphabet (IPA, International Phonetic Association, 1999) that is represented in the form of distinctive features (for details, see Anderson et al. 2018 and Rubehn et al. 2024). The conversion of individual datasets to the CLDF standard is supported by dedicated Python libraries (most importantly the CLDFBench packages, Forkel and List 2020) that help to check the overall consistency of the data.
@@
From a collection of CLDF datasets, the CLICS aggregation workflow iterates over the datasets and assembles cross-linguistic colexifications for each language variety. Here, CLICS uses an efficient method that avoids comparing $n$ words in one language against $n$ words in the same language, but rather identifies colexifications from tuples, consisting of a word form and its corresponding sense (a concept set in the Concepticon catalog), with the help of hash tables (List, 2022). In other words, the method iterates over all words in a dataset only once, instead of comparing all words against each other, which would result in large computation times.
@@
Having created a large colexification network of all CLDF datasets, the CLICS workflow analyzes this data further by computing communities, that is, partitions of nodes in a graph that show more connections to each other than to other nodes outside of the partition (Newman, 2006, 8577). Communities are inferred with the help of the Infomap algorithm (Rosvall and Bergstrom, 2008) and are used to structure the web application, by allowing users to inspect either entire communities or individual subsets of the data. The methods for data aggregation and analysis are freely accessible and can be easily applied by scholars to create their analyses of subsets of the data in CLICS or by extending the CLICS collection further, as illustrated, for example, by Tjuka (2024b).
@@ 2.3 Shortcomings of CLICS
The first shortcoming relates to the data underlying CLICS. While data from 30 datasets were aggregated in Version 3.0 (Rzymski et al., 2020), many more datasets have recently been made available via the Lexibank repository (List et al., 2022). Improving the database by increasing the number of datasets is thus one of the most urgent tasks that should be addressed in an updated version.
@@
Third, CLICS 3 provided information from more than 3,000 language varieties. However, a closer look at the data showed that only a small proportion of the included languages met the requirement set by the editors of CLICS 3 to provide elicitation glosses for at least 250 concepts. For CLICS 3, the authors instead selected 30 datasets that were officially compiled from concept lists with 250 or more items. The resulting word lists for individual languages, however, were often scarce and a larger number of the languages did not meet the originally stated coverage criterion.
Fourth, CLICS 3 offered the colexification network exclusively in the form of a GML file. Although GML is a common format for the encoding of graphs (Himsolt, 2010), accepted by many software tools, including igraph (https://igraph.org, Csárdi and Nepusz, 2006), NetworkX (https://networkx.org/, Hagberg et al., 2008), and Cytoscape (http://cytoscape.org/, Smoot et al., 2011), the format is not well-suited to share the extensive data on colexification patterns computed by CLICS 3. As a result, more transparent data formats for handling colexification data and colexification networks are needed to represent the results of the CLICS workflow in detail.
@@ 3 Materials and Methods
In the following, we will introduce all necessary steps that lead to the creation of our modified CLICS 4 database. We followed the established workflow for data aggregation used in CLICS 3 to some extent (§ 2.2). However, we present a drastic increase of data based on standardized datasets (§ 3.1), introduce an improved handling of concepts during data aggregation (§ 3.2), refine the selection of languages and concepts (§ 3.3), and make the representation of the colexification data more transparent (§ 3.4).
@@ 3.1 Data Basis
CLICS 3 was based on 30 datasets available in CLDF. Many more datasets have since been published as part of the Lexibank repository, which was first published in 2022 (List et al., 2022) as Lexibank 1 and curates data from 100 different datasets of different sizes. Of those 52 Lexibank datasets were suitable for inclusion in CLICS, because they were based on concept lists that contain 250 or more items (this criterion was used to build CLICS 3, Rzymski et al. 2020). The newest version, Lexibank 2, offers data for 134 different datasets that are all phonetically transcribed (Blum et al., 2025). For our enhanced version of CLICS, we identified 95 suitable datasets. These datasets are listed in the supplementary material accompanying this study.
@@ 3.2 Concept Handling
The colexifications in CLICS result from comparing words mapped to the standardized concept sets in Concepticon (List et al., 2016; Tjuka et al., 2023). The consequent mapping of the elicitation glosses in individual datasets to the Concepticon reference catalog has been one of the most important factors that allowed for the growth of CLICS: Version 1.0 (List et al., 2014) containing 221 language varieties and 1,280 concepts, Version 2.0 (List et al., 2018) containing 1,220 language varieties and 2,487 concepts, and Version 3.0 (Rzymski et al., 2020) containing 3,156 language varieties and 2,906 concepts. However, through the mapping of the datasets to the Concepticon, a bias for a certain number of concepts that exhibit hierarchical relations to other concepts was introduced.
@@
Already with its first launch (List et al., 2016), Conception has allowed for the definition of broad concepts that are expressed as such only in specific languages or specific linguistic areas. As an example, consider the concept sets arm or hand and foot or leg. These concept sets are expressed by individual word forms in languages such as Vietnamese tay, referring to ‘arm’ or ‘hand’, or Russian noga, referring to ‘foot’ or ‘leg’. However, many languages distinguish them further, using individual words for arm, hand, foot, and leg, respectively.
@@
Some lists in Concepticon have a linguistic area or language family as a target. Thus, the introduction of underspecified concept sets, such as ARM OR HAND or FOOT OR LEG was important, because linguists reporting on Slavic languages or particular languages in South-East Asia do not elicit both arm and hand, if they know that these are always colexified in the languages under study. However, this kind of lexical underspecification, as we encounter it in the lexicons of Vietnamese and Russian, is one of the typical reasons for colexifications. Therefore, it is important to list such cases as true colexifications of arm and hand, as well as foot and leg. The original aggregation technique used by CLICS ignores these cases. As a result, important colexification information for a large number of languages is lost.
@@
In our updated version CLICS 4, we account for underspecification directly, by defining a list of 85 concept sets that exhibit underspecification along with the more specific target concepts that they cover. While most of these underspecified concept sets can be represented by two concept sets, some are represented by more than two (specifically kinship terms like sister, which has four counterparts: younger sister (of man), younger sister (of woman), older sister (of man), and older sister (of woman)). In addition, we decided to replace some concept sets with a too broad or too narrow definition by more common concept sets (e.g. replacing stone or rock by stone because rock did not occur in the data).
@@
When encountering words that are mapped to these concepts during the initial iteration over all word lists in the data, the respective words are multiplied and each of the words is mapped to the specific concept sets covered by the underspecified concept sets. Word forms that are artificially multiplied in this form are marked in the resulting dataset by providing information on the original concept set. In total, we identify 85 underspecified concept sets in Concepticon that are relevant for the data in our modified version of CLICS. Of the 1,445,845 word forms in CLICS 4, 107,921 word forms result from this refinement procedure. A detailed list of the concept replacements can be found in Appendix A.